At what energy level does the Bohr hydrogen atom have diameter 8.57 nm ?
Answer #1

Radius of the atom is r = n2 * h2 * Є0/π * m * e2 ….( 1 )

it is given that diameter D = 8.57 nm = 8.57 * 10-9 m

So radius is r = D/2 = 8.57 * 10-9 m/2 = 4.285 * 10-9 m

Here we can find the energy level simply by substituting value of r in the equation ( 1 ) which gives n = 9

But we can also calculate it by calculating the energy of the energy level

Energy of hydrogen atom in an energy level n is given by

                    E = - m * e4/8 * n2 * h2 * Є02 ………. ( 2)       

From ( 1 ) and ( 2 ) we can observe that

                   E = ( - e2/8 π * Є0 ) * ( π * m * e2/n2 * h2 * Є0 )

                      = ( - e2/8 π Є0 * r) ……… ( 3 )

                E = - ( 1.6 * 10-19 )2 / ( 8 * 3.14 * 8.85 * 10-12 * 4.285 * 10-9 )

                   = - ( 2.56 * 10-38/952.60 * 10-21 )

                   = - 0.00269 * 10-17 J

                   = - ( 0.00269 * 10-17/ 1.6 * 10-19 ) eV

                  = - 0.168 eV

Now energy of an hydrogen is given by E = - 13.6 eV/n2

             So           - 0.168 = - 13.6 /n2

               Then            n2 = 13.6/0.168 = 80.9 = 81 ( appr )

                Finally        n = 9

So the hydrogen atom is in the energy level n = 9